5) 25% Question 6 (1 point) Googles stock (GOOG) was tracked for 10 days, showing closing prices of 1,123.52;1.099.6 1,090.23; 1,092.45; 1,103.57; 1,108.71; 1,120.82; 1,118.04; 1,109.68,1,121.34.Calculate the coefficient of variation of the dataset. 1) 0.011 2) -1,096.645 3) 1,120.947 4) 24.3012 5) 13,472.52 Question 7 (1 point) rrature in Grand Haven in August (fahrenheit) for 10 days was 768 79.1: 862 697 IRR.2. Calculate the coefficient of variation of the dataset
Answer #1

For the direct answer, the coefficient of variance should be between -1 to 1

so that option 1 will be correct

Mean(µ) = (1123.52 + 1099.6 + 1090.23 + 1092.45 + 1103.57 + 1108.71 + 1120.82 + 1118.04 + 1109.68 + 1121.34)/10
Mean = 11087.96/10
µ = 1108.7959

1 n -1 1
= √( (1/10-1) * (1123.52-1108.7959)2+(1099.6-1108.7959)2+(1090.23-1108.7959)2+(1092.45-1108.7959)2+(1103.57-1108.7959)2+(1108.71-1108.7959)2+(1120.82-1108.7959)2+(1118.04-1108.7959)2+(1109.68-1108.7959)2+(1121.34-1108.7959)2)
= √( (1/9) * (14.72412 + -9.19592 + -18.56592 + -16.34592+ -5.22592 + -0.08592 + 12.02412 + 9.24412 + 0.88412 + 12.54412))
= √( (1/9) * (216.79912081 + 84.56457681 + 344.69264281 + 267.18844681 + 27.31003081 + 0.00737881000001 + 144.57898081 + 85.45338481 + 0.78163281 + 157.35444481))
= √ 147.63465025
σ= 12.1505

Coefficient of Variance = σ/µ
= 12.1505 / 1108.7959
Coefficient of Variance = 0.0109=0.011